OpenStudy (anonymous):

How to create a quadratic equation when given is that one of the solutions must be 1/(3+2sqrt2) . Show steps.

OpenStudy (king):

hw many times will u ask the same question?stop spamming!

OpenStudy (anonymous):

Spamming? I didn't get an answer last time.

OpenStudy (king):

so wait u will get an answer..dont post the same question again

OpenStudy (anonymous):

y = (x- 1/(3+2sqrt2))(x-a) a can be anything

OpenStudy (anonymous):

Thanks!

OpenStudy (dumbcow):

you could also rationalize it \[\frac{1}{3+2\sqrt{2}}*\frac{3-2\sqrt{2}}{3-2\sqrt{2}} = 3-2\sqrt{2}\] so then \[x = 3\pm 2\sqrt{2}\] \[x-3 = \pm2\sqrt{2}\] \[(x-3)^{2} = 8\] \[x^{2} -6x+1 = 0\]

OpenStudy (anonymous):

Thank you!